OpenStudy (anonymous):

PLEASE HELP MEEE!! Are these functions even or odd and how did u find that? 2/x^2, 4 square root of x, x^5-2x^3+x

OpenStudy (campbell_st):

well just subsitute -x for x and see if you get the original function if f(-x) = f(x) then the function is even f(-x) = -f(x) the function is odd and lastly if neither is true, then the function is neither odd nor even thats the theory... so for the 1st \[f(-x) = \frac{2}{(-x)^2}\] can you simplify this..?

OpenStudy (anonymous):

Wouldn't the (-x)^2 just become an x^2 again? so the function is even?